# Spectrum Sharing in Satellite-Terrestrial Integrated Networks: Frameworks, Approaches, and Opportunities

### v1

@@ III-D Dynamic Spectrum Access-Based Approach
In STINs, users can sense and access the available spectrum as long as they are not causing significant interference to other users or the spectrum is idle, as shown in Fig. 3(d). Dynamic spectrum access (DSA) could be used to achieve this objective. Since the spectrum resource is scarce in STINs, DSA is a widely adopted SS approach that takes advantage of the temporal holes of the spectrum, and the spectrum efficiency and system rate can thus be significantly promoted. Since NTNs have high mobility, STIN users can adopt DSA strategies to access the NTN spectrum by integrating artificial intelligence (AI) tools. Similarly, NTNs can sense the spectrum licensed to TN users and adopt DSA to provide services for users. Hence, the DSA protocol is essential for the DSA-based approach in SS and has been designed to guarantee the performance of STINs in existing studies. The NTN and TN nodes can exchange spectrum information using a dedicated SS approach.
@@ III-E Game Theory-Based Approach
Unlike centralized control, users tend to make their own decisions and compete with each other to access the shared spectrum. Generally, all users need to interact with each other to decide the optimal SS pattern. Recognized as a powerful theoretical approach, game theory models the interacting users as rational players aiming to maximize their own benefits by selling and buying spectrum resources, as shown in Fig. 3(e). Hence, the spectrum allocation and access scheme of the TN/NTN nodes and the users can be described by adopting precise game models, and the optimal spectrum utilization pattern can be obtained by deriving the Nash equilibrium. It can be noted that, since STINs are heterogeneous with multiple hierarchies, the internal coordination of various users in SS often exhibits considerably high complexity. In this sense, the game theory approach is considered one of the most competitive approaches for SS in STINs.
FIGURE 3: Illustrations of advanced approaches to SS in STINs.
@@ IV Performance Metrics of SS in STINs
IV-A User Throughput and System Capacity of STINs
@@
User throughput and system capacity refer to the amount of data that the user and the entire system can transmit and receive over a certain period, respectively. As a key performance metric, user throughput can be obtained by dividing a user’s data flow by the period, demonstrating the efficiency of STIN in continuously providing wireless services. The system capacity is the sum of user rates and a system-level performance metric, which denotes the ability of STIN to provide services in some geographical regions. Since STINs have dynamic architectures, a high and stable expected system capacity implies that STINs can provide wireless services with nice qualities.
IV-B Area Spectrum Efficiency of NTN and/or TN
Area spectrum efficiency (ASE) refers to the spectrum efficiency of a unit area. Since STINs consist of various network components with different densities, ASE can be used to evaluate the average performance of a large-scale network from a system-level perspective. For a specific network segment, the ASE can be calculated by multiplying the density of the network components by the average spectrum efficiency of the network. By determining the ASE of STINs, we can also assess their system capacity per unit of bandwidth.
IV-C Interference Intensity between NTN and TN
@@
Interference intensity refers to the power strength per unit of bandwidth caused by secondary users to primary users, which denotes the negative impact of secondary users when accessing the shared spectrum in STINs. In SS, the users receive interference from the TN nodes during NTN downlink transmission, and the TN nodes receive interference from the users during NTN uplink transmission. Hence, a smaller interference intensity implies a higher efficiency and capacity of the STINs. Since the total interference power grows as the bandwidth increases, the STIN performance can be guaranteed by maintaining a satisfactory interference intensity.
IV-D Spectrum Access Probability of NTN User
@@
Spectrum access probability (SAP) refers to the likelihood that a secondary user will access the shared spectrum when they want to transmit packets. This metric indicates how fairly the spectrum is utilized among secondary users. When a channel is detected as busy, the SAP is low, prompting secondary users to either attempt to re-access the spectrum later or to utilize their reserved spectrum. For secondary users, such as NTN users or satellites within STINs, we can analyze or simulate the SAP when they employ an SS approach. Note that SAP is typically included in the average data rate calculations.
IV-E End-to-End Latency in STINs
@@
End-to-end (E2E) latency refers to the time a user receives a data packet across the network once the packet is triggered. This metric is essential in the SS of STINs, as it reflects the timeliness and fairness of packet transmission. In STINs, secondary users may need to delay their transmissions if they cannot access the spectrum in the current time slot. Additionally, when dealing with satellite-to-ground communication links, it is essential to consider propagation delay, which can add several milliseconds to the E2E latency. Generally, a smaller E2E latency indicates a quicker packet transmission.
IV-F Energy Efficiency of NTN and/or TN
@@
Energy efficiency (EE) refers to the number of bits a user can transmit by consuming per unit of power. As a key performance metric, EE is obtained by dividing data throughput by the amount of power consumption, which denotes energy utilization in transmitting signals. Since NTN nodes and users in STINs often face limited energy supply restrictions, it is crucial to promote their EE to transmit more data and provide better QoS. Meanwhile, increasing EE also enables us to better utilize scarce spectrum resources. Thus, a good guarantee of EE is essential for the overall system performance of STINs.
@@ V Case Study: SS in STINs Using Protection Zone-based Approach
In this section, we quantitatively study the effectiveness of SS in STINs by adopting the protection zone-based approach and compare it with STINs without SS. The simulation scenarios are first depicted in Fig. 4. The NTN downlink and TN downlink framework is considered, where the NTN users are the primary users and the TN users are the secondary users. A protection zone-based approach is adopted, so that the BSs are located at least $R_{\text{p}}$ away from NTN users. For the case of SS in Fig. 4(a), NTN users are primary users each having a protection zone with radius $R_{\text{p}}$. An NTN user will use the reserved spectrum if there is a BS inside the protection zone and will use the shared spectrum if there is no BS inside. If SS is not adopted in STINs as shown in Fig. 4(b), NTN users and TN users will experience interference from the NTN spectrum and TN spectrum, respectively.
The simulation results are shown in Fig. 5. Both the system capacity and the average data rate of STINs without SS remain almost the same for different reserved NTN bandwidths $B_{\text{R}}$ in STINs with SS. Specifically, as shown in Fig. 5(a), the capacity of TN with SS first surpasses that without SS when $50\leq B_{\text{R}}\leq 110$, but is exceeded by the latter subsequently. The capacity of NTN with SS is always higher than that without SS, and the sum capacity of STINs with SS continues to increase. In Fig. 5(b), the average data rate of the users in STINs with SS is compared to that of the users in STINs without SS. With an increase of $B_{\text{R}}$, the data rate of both NTN users and TN users using shared spectrum decreases because more bandwidth is allocated to the reserved spectrum. However, the data rate of the NTN user using reserved spectrum increases. Therefore, it can be concluded that a win-win solution can be obtained by using SS within a certain range of reserved NTN bandwidth.
@@
FIGURE 5: System capacity and average data rate of STINs with SS and without SS. The total bandwidth is $300$ MHz, while the NTN bandwidth and the TN bandwidth are fixed to $120$ MHz and $180$ MHz, respectively, for STINs without SS. The BSs are randomly distributed in a $200$ km×$\times 200$ km remote area with a density of $\lambda_{\text{T}}=3\times 10^{-3}$/km2. Each BS has a service radius of $R_{\text{T}}=10$ km and transmit power of $P_{\text{T}}=46$ dBm. The pathloss exponent in TNs is $\alpha_{\text{T}}=3.5$. In NTNs, the satellites are distributed on an orbital sphere with a density of $\lambda_{\text{N}}=1\times 10^{-5}$/km2 at an example altitude of $H_{\text{N}}=500$ km. The transmit power, main-lobe gain, side-lobe gain, and pathloss exponent are $P_{\text{N}}=50$ dBm, $G_{\text{ml}}=30$ dBi, $G_{\text{sl}}=10$ dBi, and $\alpha_{\text{N}}\approx 2.0$, respectively. The noise power at the user side is $-110$ dBm.
@@ VI Future Opportunities of SS in STINs
In this section, we briefly discuss future applications of SS in STINs. An overview of these potential applications is presented in Fig. 6, and detailed descriptions of each of them are then given as follows.
@@ VI-B SS for Multi-Connectivity-Based Cooperative STINs
Multi-connectivity in STINs indicates that a user can simultaneously connect to multiple network nodes, including BSs or access points (APs) in TNs and satellites in NTNs. Considering the complexity of the environment, cooperative communication, which includes relaying, joint transmission or reception, etc., can be utilized to connect different network tiers in STINs and promote more resilient and adaptable communications. However, the bandwidth allocated to a user, which may have multiple connection links, can be restricted due to the scarcity of available spectrum resources. Hence, allocating spectrum resources among multiple users, each possessing multiple links, presents significant research opportunities. In addition, the organization of cooperative network nodes, whether centralized or decentralized, requires further exploration and investigation.
VI-C Unlicensed SS in STINs
Unlicensed frequency bands offer a significant amount of high-bandwidth and low-frequency spectrum resources. However, users operating in these unlicensed bands, such as Wi-Fi, typically use the carrier sense multiple access with collision detection (CSMA/CD) medium access control (MAC) protocol to communicate in close proximity. As a result, these unlicensed bands are often underutilized in remote areas. To address this issue, satellite communications can opportunistically access unlicensed bands to maximize the use of available spectrum. It is essential to design modified MAC protocols that effectively achieve unlicensed SS in STINs. Additionally, the physical layer must adapt its frame structure to accommodate satellite mobility, the varying fading channels encountered in space and space-to-ground links, and the challenges of synchronization due to the extended propagation delays between satellites and ground users.
VI-D Multiple Access-Enabled Cognitive SS in STINs
@@
Multiple access techniques can be incorporated into STINs due to their shared goals of improving spectrum utilization efficiency. Multiple access allows multiple users to be served simultaneously in the same frequency band. Using this characteristic, multiple access and cognitive SS can be connected by sensing the real-time radio environment. The cognitive radio can sense the surroundings and adjust the transmission parameters accordingly. With these sensing results, secondary users can assess the unutilized spectrum without interfering with primary users. In addition, some other benefits come with this fusion as well, including increased network capacity, more efficient and flexible resource allocation, and decreased interference. By continuously monitoring the radio environment and adjusting transmission parameters, more reliable network performance can be ensured, and a more equitable distribution of spectrum resources among users can be promoted.
VI-E SS for Different Satellite Constellations
@@
As satellite constellations have expanded to mega size in recent years, the number of different satellite constellations built by different satellite companies is also increasing. Satellite constellations can also share their proprietary spectrum resources to increase the system’s capacity. Due to the mobile nature of the satellites, the NTN architecture is always dynamic. It changes rapidly, leading to the significant challenge of maintaining reliable and high-quality services under the restriction of limited and congested spectrum resources. Different satellite constellations face their transmission conditions as service providers. They require compatible spectrum access strategies. SS techniques enable the high-efficiency utilization of spectrum resources while introducing controllable interference. Therefore, identifying efficient SS configurations is crucial for various satellite constellations to mitigate interference and maintain service quality. Meanwhile, to meet the constellation dynamics, the designed SS strategy should also be adaptive to available satellites and spectrum resources from time to time in providing wireless services.
VI-F SS for Space-Air-Ground Integrated Networks (SAGINs)
Due to the rapid growth of aircraft and unmanned aerial vehicles (UAVs), STINs can be significantly extended by incorporating these users as the aerial network nodes to evolve towards the architecture of space-air-ground integrated networks (SAGINs). Compared to satellites, aerial nodes are more flexible in SAGIN architectures and much closer to the users. Therefore, SS is also a critical issue for aerial nodes, as these nodes can promptly monitor the spectrum dynamics, detect spatial and temporal holes, and provide corresponding wireless services for SAGIN users under well-designed SS strategies. Meanwhile, aerial nodes can act as interims between space and ground by forwarding messages, offloading, and balancing tasks among different regions. These tasks also require SS strategy designs to coordinate the utilization of spectrum resources among various SAGIN users.
@@ VII Conclusions
Chungang Yang (guideyang2050@163.com) Chungang Yang is a full professor at Xidian University, leading the GUIDE: Game, Utility, Artificial Intelligent Design for Emerging Communications research team. His research interests include artificial intelligence 6G wireless mobile networks, intent-driven networks, space-terrestrial networks, and game theory for emerging communication networks.
Chao Ren (chaoren@ustb.edu) Chao Ren received a Ph.D. from Xidian University, China, in 2017. He is an Associate Professor at the University of Science and Technology in Beijing, China. He served as a Review Editor for Frontiers in Computer Science, the Co-Chairs for Ucom23-24 WKSP, a Guest Editor for Space-Integrated-Ground Information Network, and a TPC Member for IEEE ICC and Globecom.
Haijun Zhang (haijunzhang@ieee.org) Haijun Zhang is a Full Professor and Vice Dean at the University of Science and Technology in Beijing, China. He serves/served as an Editor of IEEE Transactions on Information Forensics and Security, IEEE Transactions on Communications, IEEE Transactions on Network Science and Engineering, and IEEE Transactions on Vehicular Technology. He received the IEEE CSIM Technical Committee Best Journal Paper Award in 2018, the IEEE ComSoc Young Author Best Paper Award in 2017, and the IEEE ComSoc Asia-Pacific Best Young Researcher Award in 2019. He is a Fellow of IEEE.

### v2

@@ III-D Dynamic Spectrum Access-Based Approach
Method: Dynamic spectrum access (DSA) is a widely adopted SS approach that takes advantage of the temporal holes of the licensed spectrum. In STINs, users can sense and access the available spectrum as long as they are not causing significant interference to other users or the spectrum is idle, as shown in Fig. 3(d), and DSA can help achieve this objective by allowing secondary user access using the spectrum holes. In this way, the spectrum efficiency and system rate can thus be significantly promoted.
Applications: Since NTNs have high mobility, STIN users can adopt DSA strategies to access the NTN spectrum by integrating artificial intelligence (AI) tools. Similarly, NTNs can sense the spectrum licensed to TN users and adopt DSA to provide services for NTN users. Hence, the DSA protocol is essential for the DSA-based approach in SS and has been designed to guarantee the performance of STINs in existing studies.
Difficulties and Solutions: Since the NTN nodes need to sense and detect spectrum holes at various geolocations over time to provide wireless access, machine learning techniques incorporating dedicated geometric models are necessary for DSA in STINs. Meanwhile, since NTN nodes are independent, distributed NTN sensing schemes should be designed as well.
@@ III-E Game Theory-Based Approach
Method: Game theory is a powerful theoretical approach that models the interacting users as rational players aiming to maximize their own benefits by selling and buying spectrum resources, as shown in Fig. 3(e). In STINs, users tend to make their decisions and compete with each other to access the shared spectrum, and ultimately decide the optimal SS pattern by obtaining the game Nash equilibrium.
Applications: In STINs, the spectrum allocation and access scheme of the TN/NTN nodes and the users can be described by adopting precise game models, and the optimal spectrum utilization pattern can be obtained. This is particularly suitable for NTN downlink scenarios, as potential users may access the shared spectrum when deciding their access mode.
Difficulties and Solutions: Since STINs are heterogeneous with multiple hierarchies, the internal coordination of various users and network devices in SS often exhibits considerably high complexity. In this sense, the selection of appropriate game theory approach is the key issue in modeling the sophisticated interaction behaviors of STINs.
@@ IV Performance Metrics of SS in STINs
IV-A User Throughput and Area Spectrum Efficiency in STINs
@@
User throughput refers to the amount of data that the user can transmit and receive over a certain period, and system capacity is the sum of all user throughput. They both demonstrate the efficiency of STIN in continuously providing wireless service. Area spectrum efficiency (ASE) refers to the spectrum efficiency of a unit area, which can be used to evaluate the average performance of a large-scale network from a system-level perspective. Since STINs have dynamic architectures, a high and stable expected user throughput and ASE imply that STINs can provide high-quality wireless services, especially under large bandwidth and the assistance of SS.
IV-B Interference Intensity between NTN and TN
@@
Interference intensity refers to the power strength per unit of bandwidth caused by secondary users to primary users, which denotes the negative impact of secondary users when accessing the shared spectrum in STINs. SS brings additional interference into the system, as users receive interference from the TN nodes during NTN downlink transmission, and the TN nodes receive interference from the users during NTN uplink transmission. Hence, a smaller interference intensity implies a higher efficiency and capacity of the STINs. Since the total interference power grows as bandwidth increases, the STIN performance can be guaranteed by maintaining a satisfactory interference intensity.
IV-C Spectrum Access Probability of NTN User
@@
Spectrum access probability (SAP) refers to the likelihood that a secondary user will access the shared spectrum when they want to transmit packets. This metric indicates how fairly the spectrum is utilized among secondary users. When a channel is detected as busy, the SAP is low, prompting secondary users to either attempt to re-access the spectrum later or to utilize their reserved spectrum. For secondary users, such as NTN users or satellites within STINs, we can analyze or simulate the SAP when they employ an SS approach. Generally, a higher bandwidth is beneficial for SAP, as NTN users have more spectrum resources for SS.
IV-D End-to-End Latency in STINs
@@
End-to-end (E2E) latency refers to the time a user receives a data packet across the network once the packet is triggered. This metric is essential in the SS of STINs, as it reflects the timeliness and fairness of packet transmission. In STINs, secondary users may need to delay their transmissions if they cannot access the spectrum in the current time slot. Additionally, when dealing with satellite-to-ground communication links, it is essential to consider propagation delay, which can add several milliseconds to the E2E latency. Generally, a larger bandwidth indicates a higher data rate and eventually results in a smaller E2E latency.
IV-E Energy Efficiency of NTN and/or TN
@@
Energy efficiency (EE) refers to the number of bits a user can transmit by consuming per unit of power. As a key performance metric, EE is obtained by dividing data throughput by the amount of power consumption, which denotes energy utilization in transmitting signals. Since NTN nodes and users in STINs often face limited energy supply restrictions, it is crucial to promote their EE to transmit more data and provide better QoS. Meanwhile, increasing EE also enables us to better utilize scarce spectrum resources. Thus, a good guarantee of EE is essential for the overall system performance of STINs, and SS technique can naturally promote EE especially under large bandwidth conditions.
@@ V Case Study: SS in STINs Using Protection Zone-based Approach
FIGURE 5: System capacity and average data rate of STINs with SS and without SS. The total bandwidth is $300$ MHz, while the NTN bandwidth and the TN bandwidth are fixed to $120$ MHz and $180$ MHz, respectively, for STINs without SS. The BSs are randomly distributed in a $200$ km×$\times 200$ km remote area with a density of $\lambda_{\text{T}}=3\times 10^{-3}$/km2. Each BS has a service radius of $R_{\text{T}}=10$ km and transmit power of $P_{\text{T}}=46$ dBm. The pathloss exponent in TNs is $\alpha_{\text{T}}=3.5$. A typical TN user is assumed to be located at the center of the serving area, while typical NTN users are $50$ km away from the typical TN user. The noise power at the user side is $-110$ dBm. In NTNs, the satellites are distributed on an orbital sphere with a density of $\lambda_{\text{N}}=1\times 10^{-5}$/km2 at an altitude of $500$ km. The transmit power, main-lobe gain, side-lobe gain, and path loss exponent are $P_{\text{N}}=50$ dBm, $G_{\text{ml}}=30$ dBi, $G_{\text{sl}}=10$ dBi, and $\alpha_{\text{N}}\approx 2.0$, respectively.
In this section, we quantitatively study the effectiveness of SS in STINs by adopting the protection zone-based approach. The simulation scenarios are first depicted in Fig. 4. The NTN downlink and TN downlink framework is considered, where the NTN users are the primary users and the TN users are the secondary users. A protection zone-based approach is adopted, so that the BSs are located at least $R_{\text{p}}$ away from NTN users. For the case of SS in Fig. 4(a), NTN users are primary users each having a protection zone with radius $R_{\text{p}}$. An NTN user will use the reserved spectrum if there is a BS inside the protection zone and will use the shared spectrum if there is no BS inside. If SS is not adopted in STINs as shown in Fig. 4(b), NTN users and TN users will experience interference from the NTN spectrum and TN spectrum, respectively. This case study evaluates the aforementioned networks using two key performance metrics. First, the average data rate is determined by multiplying the product of spectral efficiency and bandwidth. Second, the system capacity of the NTN (or TN) is assessed by multiplying the spectral efficiency by the satellite (or BS) density, the surface area of satellites (or BSs), and the available bandwidth.
The simulation results are shown in Fig. 5, comparing the following three scenarios, i.e., Scenario 1: SS with the protection zone-based approach; Scenario 2: SS without the protection zone-based approach; Scenario 3: the setting without SS. Note that as reserved NTN bandwidth is considered in Scenario 2 and Scenario 3, the system capacity and average data rate of them remain almost the same. Specifically, in Fig. 5(a), the TN system capacity in Scenario 1 first surpasses that of Scenario 2 when $B_{\text{R}}$ is around $100$ MHz, but the latter subsequently exceeds it. After applying SS in Scenario 2, the excessive bandwidth can be utilized more efficiently, leading to an increase in sum and NTN system capacity compared to Scenario 3. The protection zone-based approach in Scenario 1 can utilize excessive reserved bandwidth and decrease co-frequency interference by setting an empirical protection radius $R_{\text{p}}$. If a BS is located within distance $R_{\text{p}}$, the NTN user will switch to the shared spectrum instead of the reserved spectrum. A gradual increase in both the sum and NTN system capacity can be observed.
In Fig. 5(b), the average data rate of the users in STINs with SS is compared to that of the users in STINs without SS. Compared to Scenario 3, there will be higher data rate for NTN user but lower data rate to TN user in Scenario 2, which is mainly due to stronger co-frequency interference from satellite networks. In Scenario 1, with an increase of $B_{\text{R}}$, the data rate of the NTN user and the TN user using shared spectrum decreases because more bandwidth is allocated to the reserved spectrum. However, the data rate of the NTN user using reserved spectrum increases. Therefore, by selecting the appropriate reserved bandwidth according to implementation requirements, a win-win solution can be obtained by using SS with the protection zone design.
@@ VI Future Opportunities of SS in STINs
In this section, we briefly discuss future applications of SS in STINs. An overview of these potential applications is presented in Fig. 6, and detailed descriptions of each of them are given as follows.
@@ VI-B SS for Multi-Connectivity-Based Cooperative STINs
Multi-connectivity in STINs, which has been referred to in the 3GPP Technical Standard [14], indicates that a user can simultaneously connect to multiple network nodes, including BSs or access points in TNs and satellites in NTNs. With proper synchronization and mitigation of delay, phase misalignment, and Doppler shifts with algorithms [15], cooperative communication, e.g., relaying, joint transmission, or reception, can be utilized to connect different network tiers in STINs and promote more resilient and adaptable communications. However, the bandwidth allocated to a user, which may have multiple connection links, can be restricted due to the scarcity of available spectrum resources. Hence, the allocation of spectrum resources among multiple users, each possessing multiple links, presents significant research opportunities. In addition, the organization of cooperative network nodes, whether centralized or decentralized, requires further exploration and investigation.
VI-C Multiple Access-Enabled Cognitive SS in STINs
@@
Multiple access allows multiple users to be served simultaneously in the same frequency band. Using this characteristic, multiple access and cognitive SS can be connected by sensing the real-time radio environment. The cognitive radio can sense the surroundings and adjust the transmission parameters accordingly. With these sensing results, secondary users can assess the unutilized spectrum without interfering with primary users. In addition, this fusion also offers some other benefits, including increased network capacity, more efficient and flexible resource allocation, and decreased interference. By continuously monitoring the radio environment and adjusting transmission parameters, more reliable network performance can be ensured, and a more equitable distribution of spectrum resources among users can be promoted.
VI-D SS for Different Satellite Constellations
@@
As satellite constellations have expanded to mega size in recent years, the number of different satellite constellations built by different satellite companies is also increasing. Satellite constellations can also share their proprietary spectrum resources to improve the system’s capacity. Due to the mobile nature of the satellites, the NTN architecture is always dynamic. It changes rapidly, leading to the significant challenge of maintaining reliable and high-quality services under the restriction of limited and congested spectrum resources. Different satellite constellations face their transmission conditions as service providers. They require compatible spectrum access strategies. SS techniques enable the high-efficiency utilization of spectrum resources while introducing controllable interference. Meanwhile, to meet the constellation dynamics, the designed SS strategy should also be adaptive to available satellites and spectrum resources from time to time in providing wireless services.
@@ VII Conclusions
Chungang Yang (guideyang2050@163.com) is a full professor at Xidian University, leading the GUIDE: Game, Utility, Artificial Intelligent Design for Emerging Communications research team. His research interests include artificial intelligence 6G wireless mobile networks, intent-driven networks, space-terrestrial networks, and game theory for emerging communication networks.
Chao Ren (chaoren@ustb.edu) is an Associate Professor and Deputy Director with the University of Science and Technology in Beijing, China. He is an Editorial Board Member for the Journal of Communications and Frontiers in Computer Science, recognized with an Outstanding Editor Award. He has been a Guest Editor for Space-Integrated-Ground Information Networks and Information Security Research, Co-Chair of the UCOM 2023-2025 Workshop, and Technical Program Committee member for over 30 international conferences in communications.
Haijun Zhang (haijunzhang@ieee.org) is a full professor and dean at the University of Science and Technology in Beijing, China. He serves/served as an Editor of IEEE Transactions on Information Forensics and Security, IEEE Transactions on Communications, IEEE Transactions on Network Science and Engineering, and IEEE Transactions on Vehicular Technology. He received the IEEE CSIM Technical Committee Best Journal Paper Award in 2018, the IEEE ComSoc Young Author Best Paper Award in 2017, and the IEEE ComSoc Asia-Pacific Best Young Researcher Award in 2019. He is a Fellow of IEEE.